• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name gilman

    Log entry time 16:02:46 on June14,2000

    Entry number 43737

    keyword=day shift summary June 14, 2000

    start of shift: Q3 off, Joe M doing ep energy measurement

    9:50: called for controlled access, Ed Folts et al going in for more
    magnet work...

    13:05 Ed calls up and says we are fixed...

    13:30 Ask for beam back...

    13:55 start using beam


    time run events comments
    2035 bad prescales
    2036 C pointing run
    2037 HARP scan run with carbon in, end of run
    saved late, after target move to 4 cm started
    2038 4cm dummy pointing run
    2039

    realize spectrometer set to 53.79, should be 53.65, reset it...

    2040 15 cm dummy pointing run
    2041 carbon pointing run
    2042 4 cm dummy pointing run

    3:05-3:27: RF cavity dead, being bypassed

    2043 10 cm dummy pointing run
    set collimators to 6 msr each arm
    2044 spot check, setup for efficiency test
    15:40 2045 pion prodcuction / efficiency test at 5 ua

    don't quite make it to start of production data by end of shift,
    but Q3 looks to be fixed...